# Characterizing Hitchin representations over real closed fields

### v2

@@ 3.4. $\mathbb{F}$-Hitchin component and representations
The first statement follows from 3.7. The second statement follows since the extension of semi-algebraic sets to $\mathbb{F}$ preserves semi-algebraically connected components, see 2.5 and 2.9. ∎
@@ 5.2. Total positivity
We say that $F_{1}$ is the ascending flag and $F_{3}$ is the descending flag of $\mathcal{B}$. The following proposition shows that the double ratios mimic the behaviour of the cross-ratios for $n=2$ in higher dimensional flag varieties.
@@ 6.1. $\mathbb{F}$-Hitchin representations are positively hyperbolic
Recall from 3.11 that an $\mathbb{F}$-Hitchin representation is a representation that lies in the $\mathbb{F}$-extension $\mbox{\rm Hom}_{\textnormal{Hit}}(\pi_{1}(S),\mbox{\rm PSL}(n,\mathbb{F}))$ of the set of real Hitchin representations $\mbox{\rm Hom}_{\textnormal{Hit}}(\pi_{1}(S),\mbox{\rm PSL}(n,\mathbb{R}))$.
@@ 6.2. $\mathbb{F}$-positive representations are positively hyperbolic
In the following we prove that $\mathbb{F}$-positive representations (Definition 1.4) are positively hyperbolic and dynamics preserving in the following sense.
@@
An $\mathbb{F}$-positive representation $\rho\colon\pi_{1}(S)\to\mbox{\rm PSL}(n,\mathbb{F})$ with $\rho$-equivariant limit map $\xi_{\rho}$ is dynamics-preserving if for all $e\neq\gamma\in\pi_{1}(S)$ we have $\xi_{\rho}(\gamma^{+})=F_{\rho(\gamma)}^{+}$, where $\gamma^{+}$ denotes the attracting fixed point of $\gamma$ and $F_{\rho(\gamma)}^{+}$ the stable flag of $\rho(\gamma)$; compare 5.11.
This result will be crucial in proving that $\mathbb{F}$-positive representations are $\mathbb{F}$-Hitchin. In the case $\mathbb{F}=\mathbb{R}$, this result is proven implicitly in [FG06, Theorem 7.2]. However they first extend the limit map continuously to a positive map on the whole $\partial_{\infty}\tilde{S}$, and then use continuity and dynamical arguments to prove the statement. We provide a proof that uses positivity of the limit map directly.
@@
Let $\mathbb{F}$ be a real closed field and $\rho\colon\pi_{1}(S)\to\mbox{\rm PSL}(n,\mathbb{F})$ an $\mathbb{F}$-positive representation. Then $\rho$ is positively hyperbolic and dynamics-preserving. In particular, the limit map of an $\mathbb{F}$-positive representation is unique.
@@ 7. A variant of the Bonahon–Dreyer coordinates
In this section we describe an explicit semi-algebraic model for the Hitchin component in the sense of 3.5. It is based on the seminal work of Fock–Goncharov [FG06]. Originally defined for non-closed surfaces, Bonahon–Dreyer [BD14] adapt the coordinates to closed surfaces. In the case $n=2$ they agree with the shear coordinates of Teichmüller space, see [Thu98, Section 9] or [Bon96, Theorem A]. We first focus on the case $\mathbb{F}=\mathbb{R}$, and later extend to other real closed fields in Section 7.5. The coordinates build on the existence of positive limit maps into the flag variety associated to Hitchin representations.
@@ 7.3. The parametrization
7.4. The Bonahon–Dreyer coordinates as a semi-algebraic model
This parametrization defines a semi-algebraic model for the Hitchin component in the sense of 3.5.
@@ 7.5. Extension to real closed fields
By 3.7, $\textnormal{pr}^{\textnormal{BD}}_{\mathbb{F}}$ induces a bijection between $\textnormal{Hit}(S,n)_{\mathbb{F}}$ and $\mathcal{X}_{\mathbb{F}}$. ∎
8. Proof of the main theorem
@@
for all $a=1,\ldots,n-1$. This follows from the second statement in 6.4, that says that $\rho$ is dynamics-preserving, i.e. $\xi_{\rho}(\gamma^{+})=F_{\rho(\gamma)}^{+}$. Thus $\Psi$ satisfies conditions (A)-(D), in other words $\textnormal{Im}(\Psi)\subseteq\mathcal{X}_{\mathbb{F}}$. The forward direction of 1.1 implies that $\textnormal{Im}(\Psi)=\mathcal{X}_{\mathbb{F}}$. Since the triple and double ratios are ${\rm PGL}(n,\mathbb{F})$-invariant, conjugate $\mathbb{F}$-positive representations will have the same value under $\Psi$, see 7.2.

### v3

@@ 3.3. $\mathbb{F}$-Hitchin component and representations
The first statement follows from 3.8. The second statement follows since the extension of semi-algebraic sets to $\mathbb{F}$ preserves semi-algebraically connected components, see 2.5 and 2.9. ∎
@@
3.4. Real spectrum compactification of the Hitchin component
We have seen in Section 2.3 how to define the real spectrum compactification of a closed semi-algebraic set, which we now would like to apply in the context of character varieties and the Hitchin component.
Let $G$ be either $\mbox{\rm SL}(n,\mathbb{R})$ or $\mbox{\rm PSL}(n,\mathbb{R})$. In Section 3.2, we alluded to the fact that there exists a semi-algebraic model $p$ for $\chi(S,G)$, that identifies $\chi(S,G)$ with a closed semi-algebraic subset $Y(S,G)\coloneqq\textnormal{Im}(p)$ of some $\mathbb{R}^{d}$. We can embed $\chi(S,G)$ in the compact space
$\textnormal{RSp}(\chi(S,G))\coloneqq\widetilde{Y(S,G)}\subseteq\textnormal{RSp}(\mathbb{R}[X_{1},\ldots,X_{d}])$
with dense image, and hence the latter provides a compactification of $\chi(S,G)$.
To get a better description of the points in $\chi(S,G)$, we need to introduce some notation. Recall that if $\mathbb{F}$ is a real closed field, we write $G_{\mathbb{F}}$ for the $\mathbb{F}$-extension of $G$, which agrees with either $\mbox{\rm SL}(n,\mathbb{F})$ or $\mbox{\rm PSL}(n,\mathbb{F})$, see 3.3.
Definition 3.13.
Let $\mathbb{F}$ be a real closed field and $\rho\colon\pi_{1}(S)\to G$ a homomorphism. Then $\mathbb{F}$ is $\rho$-minimal if there is no proper real closed subfield $\mathbb{K}\subseteq\mathbb{F}$ such that $\rho$ is conjugate into $G_{\mathbb{K}}$ by an element of $G_{\mathbb{F}}$.
Proposition 3.14 ([BIPP23, Corollary 7.9]).
If $\rho\colon\pi_{1}(S)\to G_{\mathbb{F}}$ is reductive, then there exists a unique real closed subfield $\mathbb{F}_{\rho}\subseteq\mathbb{F}$ such that $\rho$ is $G_{\mathbb{F}}$-conjugate to a representation $\rho^{\prime}\colon\pi_{1}(S)\to G_{\mathbb{F}_{\rho}}$ and $\mathbb{F}_{\rho}$ is $\rho^{\prime}$-minimal.
Theorem 3.15 ([BIPP23, Theorem 1.1]).
Let $G$ be either $\mbox{\rm SL}(n,\mathbb{F})$ or $\mbox{\rm PSL}(n,\mathbb{F})$. Then
$\textnormal{RSp}(\chi(S,G))\cong\left\{(\rho,\mathbb{F}_{\rho})\ \middle|\begin{array}[]{l}\rho\colon\pi_{1}(S)\to G_{\mathbb{F}_{\rho}}\textrm{ reductive,}\\ \mathbb{F}_{\rho}\supseteq\mathbb{R}\textrm{ real closed and }\rho\textrm{-minimal}\end{array}\right\}_{\Big{/}\sim}\,,$
where two representations $\rho_{1}\colon\Gamma\to G_{\mathbb{F}_{\rho_{1}}}$ and $\rho_{2}\colon\Gamma\to G_{\mathbb{F}_{\rho_{2}}}$ are equivalent ($\sim$) if there exists an order-preserving isomorphism $i\colon\mathbb{F}_{\rho_{1}}\to\mathbb{F}_{\rho_{2}}$ such that $\rho_{2}$ and $i\circ\rho_{1}$ are $G_{\mathbb{F}_{\rho_{2}}}$-conjugate. Points in the boundary correspond to representations into $\mbox{\rm PSL}(n,\mathbb{F}_{\rho})$ for $\mathbb{F}_{\rho}$ a non-Archimedean field. Moreover $\mathbb{F}_{\rho}$ is of finite transcendence degree over $\mathbb{R}$.
The Hitchin component $\textnormal{Hit}(S,n)\subseteq\chi(S,\mbox{\rm PSL}(n,\mathbb{R}))$ is again closed semi-algebraic as it is a connected component of a closed semi-algebraic set. We denote by $\textnormal{RSp}(\textnormal{Hit}(S,n))$ its real spectrum compactification, i.e. the closure of $\textnormal{Hit}(S,n)$ in $\textnormal{RSp}(\chi(S,\mbox{\rm PSL}(n,\mathbb{R})))$. Recall from 1.3 that a representation $\rho\colon\pi_{1}(S)\to\mbox{\rm PSL}(n,\mathbb{F})$ is $\mathbb{F}$-Hitchin if its $\mbox{\rm PSL}(n,\mathbb{F})$-equivalence class lies in the $\mathbb{F}$-extension of $\textnormal{Hit}(S,n)$.
Theorem 3.16 ([BIPP21b, Theorem 46]).
$\textnormal{RSp}(\textnormal{Hit}(S,n))\cong\left\{(\rho,\mathbb{F}_{\rho})\ \middle|\begin{array}[]{l}\rho\colon\pi_{1}(S)\to\mbox{\rm PSL}(n,\mathbb{F}_{\rho})\textrm{ is }\mathbb{F}_{\rho}\textrm{-Hitchin,}\\ \mathbb{F}_{\rho}\supseteq\mathbb{R}\textrm{ real closed and }\rho\textrm{-minimal}\end{array}\right\}_{\Big{/}\sim}\,,$
for the same equivalence relation as in the theorem before. Furthermore, $\mathbb{F}_{\rho}=\overline{\mathbb{R}(\mbox{\rm tr}(\textnormal{Ad}(\rho)))}^{r}$, where Ad is the adjoint representation of $\mbox{\rm PSL}(n,\mathbb{F}_{\rho})$. In addition, $(\rho,\mathbb{F}_{\rho})$ represents a closed point if and only if $\mathbb{F}_{\rho}$ is Archimedean over the ring of traces $\mathbb{R}[\mbox{\rm tr}(\textnormal{Ad}(\rho))]$ of $\textnormal{Ad}\circ\rho$.
Loosely speaking, the goal of the following chapters is to replace the word “$\mathbb{F}_{\rho}$-Hitchin” by the word “$\mathbb{F}_{\rho}$-positive” in this theorem.
@@ 5.2. Total positivity
We say that $F_{1}$ is the ascending flag and $F_{3}$ is the descending flag of $\mathcal{B}$. The following proposition shows that the double ratios mimic the monotonicity behavior of the cross-ratios for $n=2$ in higher dimensional flag varieties.
@@ 5.3. Positive hyperbolicity
Note that if $\mathbb{F}\neq\mathbb{R}$, then a stable flag is not necessarily attracting in the dynamical sense.
@@ 6.1. $\mathbb{F}$-Hitchin representations are positively hyperbolic
Recall from 3.12 that an $\mathbb{F}$-Hitchin representation is a representation that lies in the $\mathbb{F}$-extension $\mbox{\rm Hom}_{\textnormal{Hit}}(\pi_{1}(S),\mbox{\rm PSL}(n,\mathbb{F}))$ of the set of real Hitchin representations $\mbox{\rm Hom}_{\textnormal{Hit}}(\pi_{1}(S),\mbox{\rm PSL}(n,\mathbb{R}))$.
@@ 6.2. $\mathbb{F}$-positive representations are positively hyperbolic
In the following we prove that $\mathbb{F}$-positive representations (Definition 1.4) are positively hyperbolic and weakly dynamics-preserving in the following sense.
@@
An $\mathbb{F}$-positive representation $\rho\colon\pi_{1}(S)\to\mbox{\rm PSL}(n,\mathbb{F})$ with $\rho$-equivariant limit map $\xi_{\rho}$ is weakly dynamics-preserving if for all $e\neq\gamma\in\pi_{1}(S)$ we have $\xi_{\rho}(\gamma^{+})=F_{\rho(\gamma)}^{+}$, where $\gamma^{+}$ denotes the attracting fixed point of $\gamma$ and $F_{\rho(\gamma)}^{+}$ the stable flag of $\rho(\gamma)$; compare 5.11.
This result will be crucial in proving that $\mathbb{F}$-positive representations are $\mathbb{F}$-Hitchin. In the case $\mathbb{F}=\mathbb{R}$, this result is proven implicitly in [FG06, Theorem 7.2]. They first extend the limit map continuously to a positive map on the whole $\partial_{\infty}\tilde{S}$, and then use continuity and dynamical arguments to prove the statement—a strategy that cannot be adapted to general real closed fields. We provide a proof that uses positivity of the limit map directly.
@@
Let $\mathbb{F}$ be a real closed field and $\rho\colon\pi_{1}(S)\to\mbox{\rm PSL}(n,\mathbb{F})$ an $\mathbb{F}$-positive representation. Then $\rho$ is positively hyperbolic and weakly dynamics-preserving. In particular, the limit map of an $\mathbb{F}$-positive representation is unique.
@@ 7. A variant of the Bonahon–Dreyer coordinates
In this section we describe an explicit semi-algebraic model for the Hitchin component in the sense of 3.6. It is based on the seminal work of Fock–Goncharov [FG06]. Originally defined for non-closed surfaces, Bonahon–Dreyer [BD14] adapt the coordinates to closed surfaces. In the case $n=2$ they agree with the shear coordinates of Teichmüller space, see [Thu22, Section 9] or [Bon96, Theorem A]. We first focus on the case $\mathbb{F}=\mathbb{R}$, and later extend to other real closed fields in Section 7.5. The coordinates build on the existence of positive limit maps into the flag variety associated to Hitchin representations.
@@ 7.3. Parametrization
7.4. Bonahon–Dreyer coordinates as a semi-algebraic model
This parametrization defines a semi-algebraic model for the Hitchin component in the sense of 3.6.
@@ 7.5. Extension to real closed fields
By 3.8, $\textnormal{pr}^{\textnormal{BD}}_{\mathbb{F}}$ induces a bijection between $\textnormal{Hit}(S,n)_{\mathbb{F}}$ and $\mathcal{X}_{\mathbb{F}}$. ∎
8. Proof of 1.1
@@
for all $a=1,\ldots,n-1$. This follows from the second statement in 6.4, that says that $\rho$ is weakly dynamics-preserving, i.e. $\xi_{\rho}(\gamma^{+})=F_{\rho(\gamma)}^{+}$. Thus $\Psi$ satisfies conditions (A)-(D), in other words $\textnormal{Im}(\Psi)\subseteq\mathcal{X}_{\mathbb{F}}$. The forward direction of 1.1 implies that $\textnormal{Im}(\Psi)=\mathcal{X}_{\mathbb{F}}$. Since the triple and double ratios are ${\rm PGL}(n,\mathbb{F})$-invariant, conjugate $\mathbb{F}$-positive representations will have the same value under $\Psi$, see 7.2.